OpenStudy (anonymous):

Write the equation of the asymptotes to the hyperbola with equation: x^2 - y^2 + 4x + 8y - 21 = 0

OpenStudy (anonymous):

I got y - 2 = (x+4) and y - 2 = -(x+4), apparently I am wrong.

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

y=2-x and y=x+6